OpenStudy (anonymous):

Will someone check my 2 questions GIVING MEDALS AND FANS. (QUESTIONS BELOW)

OpenStudy (anonymous):

what are your answers

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Can u help me with this one @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Volume of a cube: \(V =s^3\) Where s = side. Plug in what we know: \(2197 =s^3\) Now to find s, we have to cube 2197, can you do that? @starrhazel123

OpenStudy (anonymous):

yes I think soo

OpenStudy (igreen):

I mean you have to find the cube root of 2197.

OpenStudy (igreen):

Um, I guess..did you find the cube root of 2197? You can just type in google: 'cube root of 2197'

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Cube root looks like this \(^3 \! \! \! \! \! \! \sqrt{35 }\) There's a 3 beside the square root sign..and since we have to find the cube root of 2744..what's your answer gonna be?

OpenStudy (anonymous):

A I would say A

OpenStudy (igreen):

Yep, you got it.